OpenStudy (anonymous):

You know that gravity changes the potential energy of falling objects into kinetic energy. Design an experiment to investigate how distance from the ground will affect a ball’s potential energy. Write out the materials you need and the steps you will follow. Record your results and observations.

OpenStudy (anonymous):

I dont know what to do to do the experiment

OpenStudy (anonymous):

It is self-explanatory. You are using materials to compare how the distance away from the ground affects potential energy. PE = MGH (Mass * Gravity * Height) Gravity is close to constant especially near the earths surface and if you want to be consistent you would use an item with the same mass all the time example: soft ball. So you want to see how height affects potential energy and it looks like it wants you to use Kinetic Energy, KE = 1/2 mv^2 and you know velocity is the change in displacement over the change in time. You can use both equations to help you see how DISTANCE affects potential energy. Be creative.

OpenStudy (anonymous):

Also review your conservation of energy law\[PE_i + KE_i=PE_f + KE_f\] PE = Potential Energy KE = Kinetic Energy i = initial f = final

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I only need a ball because if the ball gets higher it bounces higher

OpenStudy (anonymous):

The bounce is a collision that we are not worried about but if we see how fast the ball drops in a given time we can calculate velocity because velocity is \[v = \frac{ dx }{ dt }\]

OpenStudy (anonymous):

With velocity and the mass of the ball you can calculate the kinetic energy which has to be the same amount of potential energy because